# Histopathology Slide Indexing and Search: Are We There Yet?

### v1

@@ 4 Discussion
[ht] Lung Adenocarcinoma Slide Retrieval Rank Yottixel Site Yottixel Subtype SISH Site SISH Subtype RetCCL WSI Site RetCCL Subtype 1 Brain (198.5) Lung adenocarcinoma (215.0) Liver (125.0) Lung adenocarcinoma (127.0) Lung (0.80) Lung adenocarcinoma (0.80) 2 Brain (200.0) Lung adenocarcinoma (218.5) - - Lung (0.80) Lung squamous cell carcinoma (0.77) 3 Brain (206.0) Lung squamous cell carcinoma (220.0) - - Brain (0.75) Lung squamous cell carcinoma (0.75) 4 Brain (212.5) Lung adenocarcinoma (227.0) - - Liver (0.82) Lung adenocarcinoma (0.86) 5 Brain (214.0) Lung squamous cell carcinoma (227.5) - - Liver (0.85) Lung squamous cell carcinoma (0.75)
@@
[ht] Low Grade Glioma Slide Retrieval Rank Yottixel Site Yottixel Subtype SISH Site SISH Subtype RetCCL WSI Site RetCCL Subtype 1 Brain (166.0) Glioblastoma (166.0) Brain (110.0) Glioblastoma Brain (0.80) Lower Grade Glioma (0.84) 2 Brain (167.0) Glioblastoma (167.0) Brain (116.0) Glioblastoma Brain (0.81) Lower Grade Glioma (0.80) 3 Brain (171.5) Glioblastoma (171.5) Lung (119.0) Glioblastoma Liver (0.83) Lower Grade Glioma (0.81) 4 Brain (173.5) Glioblastoma (173.5) Brain (117.0) Glioblastoma Brain (0.80) Lower Grade Glioma (0.80) 5 Brain (173.5) Glioblastoma (173.5) Brain (117.0) Glioblastoma Liver (0.83) Lower Grade Glioma (0.79)
{stable}
[ht] Hepatocellular Carcinoma Slide Retrieval Rank Yottixel Site Yottixel Subtype SISH Site SISH Subtype RetCCL WSI Site RetCCL Subtype 1 Brain (180.0) Hepatocellular carcinoma (221.0) Brain (117.0) Hepatocellular carcinoma (119.0) Liver (0.87) Hepatocellular carcinoma (0.80) 2 Brain (189.0) Hepatocellular carcinoma (222.0) Brain (128.0) Hepatocellular carcinoma (120.0) Liver (0.84) Hepatocellular carcinoma (0.84) 3 Brain (191.0) Hepatocellular carcinoma (229.0) - Hepatocellular carcinoma (128.0) Liver (0.87) Cholangiocarcinoma (0.87) 4 Brain (208.0) Hepatocellular carcinoma (231.0) - - Liver (0.84) Hepatocellular carcinoma (0.84) 5 Brain (208.0) Hepatocellular carcinoma (236.0) - - Liver (0.87) Cholangiocarcinoma (0.77)

### v2

@@ 1 $relevant\_count\leftarrow 0$ $precision\_sum\leftarrow 0$ for $i=1$ to $k$ do
We choose three internal patient cases for visual review of tissue, subtype, and patch-level search results. Slide 1 is a Lung Adenocarcinoma (LUAD) case from a patient’s partial lobectomy. Slide 2 is a Low Grade Glioma (LGG) that was retrieved by surgical biopsy. Slide 3 is from a patient with Hepatocellular Carcinoma (LIHC) who underwent a liver biopsy by fine needle aspiration.
@@
[!ht] Evaluation results of different methods on large datasets for primary site retrieval task. mMV means mean Majority Voting score and mAP means mean Average Precision. Method Dataset mMV@1 mMV@3 mMV@5 mMV@10 mAP@3 mAP@5 YOTTIXEL + KimiaNet UPENN GBM 0.706 0.735 0.735 0.794 0.752 0.765 CPTAC GBM 0.606 0.606 0.667 0.727 0.646 0.644 Yale Her2 Pos 0.548 0.559 0.538 0.570 0.627 0.624 Yale Her2 Neg 0.742 0.742 0.784 0.814 0.809 0.788 Yale Trastuzumab 0.565 0.600 0.576 0.588 0.647 0.650 SISH + DenseNet UPENN GBM 0.735 0.706 0.706 0.758 0.730 0.720 CPTAC GBM 0.706 0.697 0.697 0.710 0.725 0.729 Yale Her2 Pos 0.269 0.247 0.221 0.167 0.294 0.304 Yale Her2 Neg 0.454 0.442 0.446 0.375 0.476 0.480 Yale Trastuzumab 0.329 0.321 0.333 0.295 0.345 0.356 RetCCL UPENN GBM - - - - - - CPTAC GBM 0.588 0.594 0.750 0.846 0.728 0.736 Yale Her2 Pos 0.172 0.301 0.355 0.481 0.345 0.404 Yale Her2 Neg 0.510 0.602 0.742 0.864 0.656 0.671 Yale Trastuzumab 0.212 0.247 0.353 0.420 0.382 0.428 HSHR UPENN GBM 0.794 0.824 0.765 0.735 0.816 0.809 CPTAC GBM 0.758 0.727 0.818 0.909 0.813 0.810 Yale Her2 Pos 0.247 0.323 0.323 0.301 0.429 0.449 Yale Her2 Neg 0.571 0.592 0.622 0.663 0.632 0.612 0.447 0.447 0.494 0.529 0.581 0.569 YOTTIXEL + DenseNet Yale Trastuzumab 0.224 0.224 0.235 0.247 0.318 0.353 SISH + KimiaNet Yale Trastuzumab 0.536 0.578 0.417 0.571 0.542 0.540
@@
[!ht] Evaluation results of different methods on large datasets for subtype retrieval task. Method Dataset mMV@1 mMV@3 mMV@5 mAP@3 mAP@5 YOTTIXEL + KimiaNet UPENN GBM 0.382 0.382 0.294 0.468 0.454 CPTAC GBM 0.303 0.303 0.242 0.399 0.418 SISH + DenseNet UPENN GBM 0.706 0.656 0.677 0.725 0.725 CPTAC GBM 0.676 0.667 0.656 0.686 0.681 RetCCL UPENN GBM - - - - - CPTAC GBM 0.529 0.375 0.406 0.667 0.648 HSHR UPENN GBM 0.618 0.559 0.559 0.721 0.704 CPTAC GBM 0.758 0.727 0.758 0.806 0.785